OpenStudy (anonymous):

how will you find a slope of the line in 2/3x+3y=4 ?

OpenStudy (anonymous):

\( -\frac 29 \)

OpenStudy (anonymous):

can you show the solution ?

OpenStudy (anonymous):

I have no idea, I just looked at it and the answer reveal itself.

OpenStudy (amistre64):

ax + by = c by = -ax + c y = -a/b x + c slope: -a/b

OpenStudy (amistre64):

-2/3 ---- = -2/9 3/1